# Criteria for optimal entanglement-assisted long baseline imaging protocols

### v1

@@ VI Towards a quantum-enhanced telescopy in noisy repeater-less networks
While future quantum networks promise loss-free entanglement distribution using quantum repeaters [43] in the long-term, it is of interest to study noisy repeater-less quantum telescopy protocols, where the transmission loss for the ground-based entangled photons must be considered. Note that in this scenario the original Gottesman’s scheme will show no benefit over direct transmission, because the ancilla state consists of only one single photon and would suffer the same photon losses as the astronomical photon would. However, more complex entangled-assisted protocols that use many photons could become robust against noise, as shown in [15]. Although this work demonstrated that using an N-SPE state improved phase estimation for an ideal point source (characterized by a mutual coherence function with $|g|=1$ in Eq. 3), as we have discussed in Appendix 2, their result can not be trivially generalized to an extended source (which has mutual coherent function with $|g|<1$ in Eq. 3), which indicates that whenever the source photon is not ideal, the quantum enhancement achieved with $N$-copy SPE state will be degraded.
@@
Thus, the primary goal in a realistic quantum telescopy implementation should be to improve the number of photons that survive a lossy channel. Beyond that, the QFI-bound cannot always be saturated with LOCC, especially when the transmitted states experience loss. Therefore, an important next step is to investigate practical quantum telescopy protocols that consider the presence of loss, as expected in near-term and medium-term quantum networks.
We would like to clarify that the primary focus of this letter is to provide a systematic comparison of entanglement-assisted quantum telescopy protocols in a quantum network that operates without repeaters and quantum memories. In this context, each temporal mode emitted by the source is treated as an independent system that cannot be stored, manipulated, or jointly measured. Consequently, the ancilla photons prepared in the laboratory must be supplied individually for each temporal mode. This approach contrasts with memory-based protocols [12, 13], which are expected to become more relevant in the future when fully functional quantum memories are available.
@@ VII Conclusion
Several open questions remain, including how to better understand and design practical quantum-enhanced telescopy schemes for near-term quantum networks under imperfections such as ground-photon transmission losses and mode indistinguishability. A recent analysis for the N-SPE protocol[modak2024] assuming an idealized point source, showed a strong degradation of the phase estimation with imperfections, and the performance of our scheme under the same imperfections will be of great interest. Additionally, generalizing our approach to multi-telescope scenarios, again constrained by photon transmission losses and other imperfections, is another important open problem that will be the focus of future work.
@@ Appendix A Thermal source state and optimal quantum Fisher information
$\rho_{s}=\iint d^{2}\alpha d^{2}\beta\frac{1}{\pi^{2}\det\Gamma}\exp[-(\alpha^{*},\beta^{*})\Gamma^{-1}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}]|\alpha,\beta\rangle\langle\alpha,\beta|$ (1)
@@
$\rho_{s}^{(1)}=\frac{1}{2}[|0_{A}1_{B}\rangle\langle 0_{A}1_{B}|+|1_{A}0_{B}\rangle\langle 1_{A}0_{B}|+(g|1_{A}0_{B}\rangle\langle 0_{A}1_{B}|+g^{*}|0_{A}1_{B}\rangle\langle 1_{A}0_{B}|)]$ (3)
@@
From the quantum estimation theory, the amount of information that is contained in the state in Eq. 2 can be quantified by the corresponding quantum Fisher information[24]. Focus on the single photon term $\rho_{1}$, the symmetric logarithmic derivative of it is a solution of the differential equation and is given as:
$\frac{\hat{L}_{\mu}\rho_{s}+\rho_{s}\hat{L}_{\mu}}{2}=\frac{\partial\rho_{s}}{\partial\mu}\quad\quad\mu\in\{|g|,\theta\}\quad\quad\Longrightarrow\quad\quad\hat{L}_{|g|}=\frac{1}{1-|g|^{2}}\begin{pmatrix}-g&e^{i\theta}\\ e^{-i\theta}&-g\end{pmatrix}\quad\quad\hat{L}_{\theta}=|g|\begin{pmatrix}0&ie^{i\theta}\\ -ie^{-i\theta}&0\end{pmatrix}$ (4)
@@
$\hat{H}_{|g|}[\rho_{s}]=\textrm{Tr}[\rho_{s}\hat{L}^{2}_{|g|}]=\frac{\epsilon}{1-|g|^{2}}+O(\epsilon^{2})~{}~{}~{}~{}~{}~{}\quad\quad\quad\hat{H}_{\theta}[\rho_{s}]=\textrm{Tr}[\rho_{s}\hat{L}^{2}_{\theta}]=|g|^{2}\epsilon+O(\epsilon^{2})$ (5)
@@ Appendix B Superselection rule, reference frame, and quantum estimation under SSR
$|\psi\rangle=\sum_{n=0}c_{n}|n\rangle\quad\quad\Longrightarrow\quad\quad U(\psi)|\phi\rangle=\sum_{n=0}e^{in\phi}c_{n}|n\rangle$ (6)
@@
$|\psi\rangle=\sum_{n=0}c_{n}|n\rangle\quad\Longrightarrow\quad\mathcal{U}(\rho):=\int_{0}^{2\pi}U(\psi)|\phi\rangle\langle\phi|U(\phi)^{\dagger}=\sum_{n,n^{\prime}=0}\int_{0}^{2\pi}\frac{d\phi}{2\pi}e^{i(n-n^{\prime})\phi}c_{n}c^{*}_{n}|n^{\prime}\rangle\langle n^{\prime}|=\sum_{n=0}|c_{n}|^{2}|n\rangle\langle n|$ (7)
@@
$=\sum_{n,m,n^{\prime},m^{\prime}=0}\iint_{0}^{2\pi}\frac{d\phi_{1}}{2\pi}\frac{d\phi_{2}}{2\pi}e^{i(n-n^{\prime})\phi_{1}+i(m-m^{\prime})\phi_{2}}(P^{A}_{n}\otimes P^{B}_{m})\rho^{AB}(P^{A}_{n^{\prime}}\otimes P^{B}_{m^{\prime}})$
@@
$\mathcal{E}(\rho_{s})=(1-\epsilon)\mathcal{E}(\rho_{s}^{(0)})+\epsilon\mathcal{E}(\rho_{s}^{(1)})+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)})=(1-\epsilon)\mathbb{I}^{(0)}+\epsilon\frac{\mathbb{I}^{(1)}}{2}+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)}),$
@@
$\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{|g|}[\rho_{s}^{AB}]\quad\quad\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{\theta}[\rho_{s}^{AB}]$ (13)
@@
From the additivity and Monotonicity of quantum Fisher information, we have
$\mathbb{H}_{\mu}[\rho_{s}^{AB}]+\mathbb{H}_{\mu}[\rho_{a}^{AB}]\geq\mathbb{H}_{\mu}[\rho_{s}^{AB}\otimes\rho_{a}^{AB}]\geq\mathbb{H}_{\mu}[\mathcal{E}(\rho^{AB})]\quad\quad\mu\in\{|g|,\theta\}$ (14)
@@
$\mathcal{E}(\rho^{AB})=\sum_{n,m}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})=\bigoplus_{n,m}\sigma^{AB}_{n,m}$ (18)
@@
$\sigma^{AB}_{n,m}$ $=(1-\epsilon)|f_{n,m}|^{2}|0,n_{A}\rangle\langle 0,n_{A}|\otimes|0,m_{B}\rangle\langle 0,m_{B}|+\frac{\epsilon}{2}\left([f_{n,m-1}|^{2}|0,n_{A}\rangle\langle 0,n_{A}|\otimes|1,(m-1)_{B}\rangle\langle 1,(m-1)_{B}|\right.$ (19)
$+f_{n-1,m}f^{*}_{n,m-1}g|0,n_{A}\rangle\langle 1,(n-1)_{A}|\otimes|1,(m-1)_{B}\rangle\langle 0,m_{B}|+|f_{n-1,m}|^{2}|1,(n-1)_{A}\rangle\langle 1,(n-1)_{A}|\otimes|0,m_{B}\rangle\langle 0,m_{B}|$
$+f^{*}_{n-1,m}f_{n,m-1}g^{*}|1,(n-1)_{A}\rangle\langle 0,n_{A}|\otimes|0,m_{B}\rangle\langle 1,(m-1)_{B}|\left.\right)+O(\epsilon^{2})$ (20)
Where $|0,n_{A}\rangle$ is state with 00 photon from the ancilla and $n_{A}$ photon from the source. Since $\sigma_{n,m}^{A,B}$ are density operator defined on different orthogonal Hilbert spaces, it is not difficult to see that:
$\mathbb{H}_{\mu}[\mathcal{E}(\rho^{AB})]=\sum_{n,m=0}\mathbb{H}_{\mu}[\sigma_{n,m}^{AB}]\quad\quad\quad\mu\in\{|g|,\theta\}$ (21)
where the QFI for the individual unnormalized state is discussed below in lemma 3 and is given as:
[…]
@@
The quantum Fisher information contains in the unnormalized state $\sigma_{nm}^{AB}$ is contributed by the second term in Eq. 20, which can be rewritten as:
$\sigma^{AB\epsilon}_{n,m}=\frac{(|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2})\epsilon}{2}\begin{pmatrix}\frac{|f_{n,m-1}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}&\frac{f_{n-1,m}f^{*}_{n,m-1}g}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}g\\ \frac{f^{*}_{n-1,m}f_{n,m-1}g}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}g^{*}&\frac{|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\end{pmatrix}$ (23)
in $\{|0,n_{A}\rangle|1,(m-1)_{B}\rangle,|1,(n-1)_{A}\rangle|0,m_{B}\rangle\}$ basis. And the normalized version of it is the matrix term above, donated as $\tilde{\sigma}^{AB\epsilon}_{n,m}$ is the term inside the Bracket. Similar to the calculation in section A, one could show that the symmetric logarithmic derivative of $\tilde{\sigma}^{AB\epsilon}_{n,m}$ in estimating $|g|$ and $\theta$ is given as
$\hat{L}_{|g|}[\tilde{\sigma}^{AB\epsilon}_{n,m}]$ $=\frac{1}{1-|g|^{2}}\frac{2}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\begin{pmatrix}-|f_{n-1,m}|^{2}g&f_{n-1,m}f^{*}_{n,m-1}e^{i\theta}\\ f^{*}_{n-1,m}f_{n,m-1}e^{-i\theta}&-|f_{n,m-1}|^{2}g\end{pmatrix}$
$\hat{L}_{\theta}[\tilde{\sigma}^{AB\epsilon}_{n,m}]$ $=|g|\frac{2}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\begin{pmatrix}0&if_{n-1,m}f^{*}_{n,m-1}e^{i\theta}\\ -if^{*}_{n-1,m}f_{n,m-1}e^{-i\theta}&0\end{pmatrix}$
Therefore, we have:
$\mathbb{H}_{|g|}[\sigma_{n,m}^{AB}]$ $=\frac{(|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2})\epsilon}{2}\mathbb{H}_{|g|}[\tilde{\sigma}^{AB\epsilon}_{n,m}]=\frac{\epsilon}{1-|g|^{2}}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{H}_{\theta}[\sigma_{n,m}^{AB}]$ $=\frac{(|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2})\epsilon}{2}\mathbb{H}_{\theta}[\tilde{\sigma}^{AB\epsilon}_{n,m}]=\epsilon|g|^{2}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
@@
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n,m=1}^{n+m\leq N+1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\Longrightarrow\mathbb{h}[\rho^{AB}_{a}]=\sum_{k=1}^{N}\sum_{n=1}^{k}\frac{2|f_{n,k-n}|^{2}|f_{n-1,k-n+1}|^{2}}{|f_{n,k-n}|^{2}+|f_{n-1,k-n+1}|^{2}}$ (25)
@@
Now, we can give a non-tight upper bound $h_{k}$ by upper bounding $\sum_{n=1}^{k}\sqrt{p_{n}^{k}p_{n-1}^{k}}$ from the lemma below, which shows
$h_{k}\leq q_{k}\cos(\frac{\pi}{k+2})\quad\quad\Longrightarrow\quad\quad\mathbb{h}[\rho^{AB}_{a}]=\sum_{k=1}^{N}h_{k}\leq\sum_{k=1}^{N}q_{k}\cos(\frac{\pi}{k+2})\leq\cos(\frac{\pi}{N+2})$

### v2

@@ VI Towards a quantum-enhanced telescopy in noisy repeater-less networks
While future quantum networks promise loss-free entanglement distribution using quantum repeaters [41] in the long-term, it is of interest to study noisy repeater-less quantum telescopy protocols, where the transmission loss for the ground-based entangled photons must be considered. Note that in this scenario the original Gottesman’s scheme will show no benefit over direct transmission, because the ancilla state consists of only one single photon and would suffer the same photon losses as the astronomical photon would. However, more complex entangled-assisted protocols that use many photons could become robust against noise, as shown in [15]. Although this work demonstrated that using an N-SPE state improved phase estimation for an ideal point source (characterized by a mutual coherence function with $|g|=1$ in Eq. 3), as we have discussed in Appendix D.2, their result can not be trivially generalized to an extended source (which has mutual coherent function with $|g|<1$ in Eq. 3), which indicates that whenever the source photon is not ideal, the quantum enhancement achieved with $N$-copy SPE state will be degraded.
@@
Thus, the primary goal in a realistic quantum telescopy implementation should be to improve the number of photons that survive a lossy channel.The QFI bound consistently provides a valuable upper limit for assessing the potential effectiveness of an entangled ancilla state, even under lossy conditions. However, it is important to note that this bound may not always be saturable in such scenarios. Once an entangled ancilla state with a high QFI bound is identified, the next crucial step is to explore practical measurement schemes that can help achieve this bound, thereby paving the way for novel protocols in near-term and medium-term quantum networks.
Moreover, moving beyond the traditional passive quantum telescopy scheme—where no communication is allowed—the development of local delay-line quantum memory [39] and local vacuum beam guides [40], offering optical delays of up to $\sim 100\mu$s, enables quantum telescopy protocols with one- or two-round local operations and classical communication. This advancement, combined with theoretical exploration of more noise-resilient entangled-assisted sources, could pave the way for new protocols in near-term, repeater-less quantum networks.
@@ VII Conclusion
Several open questions remain, including how to better understand and design practical quantum-enhanced telescopy schemes for near-term quantum networks under imperfections such as ground-photon transmission losses and mode indistinguishability. A recent analysis for the N-SPE protocol[modak2024] assuming an idealized point source, showed a strong degradation of the phase estimation with imperfections, and the performance of our scheme under the same imperfections will be of great interest. Additionally, generalizing our approach to multi-telescope scenarios, again constrained by photon transmission losses and other imperfections, is another important open problem that will be the focus of future work. The new tools introduced in our paper provide a framework for studying quantum Fisher information (QFI) under superselection rule constraints and for identifying the essential resources required for quantum-enhanced telescopy protocols. These developments will facilitate the search for practical, near-term implementations of quantum telescopy in repeater-less quantum networks.
@@ Appendix A Thermal source state and optimal quantum Fisher information
$\rho_{s}=\iint d^{2}\alpha d^{2}\beta\frac{\exp[-(\alpha^{*},\beta^{*})\Gamma^{-1}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}]}{\pi^{2}\det\Gamma}|\alpha,\beta\rangle\langle\alpha,\beta|$ (1)
@@
$\rho_{s}^{(1)}=$ $\frac{1}{2}[|0_{A}1_{B}\rangle\langle 0_{A}1_{B}|+|1_{A}0_{B}\rangle\langle 1_{A}0_{B}|$ (3)
$+(g|1_{A}0_{B}\rangle\langle 0_{A}1_{B}|+g^{*}|0_{A}1_{B}\rangle\langle 1_{A}0_{B}|)]$
@@
From the quantum estimation theory, the amount of information that is contained in the state in Eq. 2 can be quantified by the corresponding quantum Fisher information[24]. Focus on the single photon term $\rho_{1}$, the symmetric logarithmic derivative of it is a solution of the differential equation $\frac{\hat{L}_{\mu}\rho_{s}+\rho_{s}\hat{L}_{\mu}}{2}=\frac{\partial\rho_{s}}{\partial\mu}$ for $\mu\in\{|g|,\theta\}$ and is given as:
$\hat{L}_{|g|}=\frac{1}{1-|g|^{2}}\begin{pmatrix}-g&e^{i\theta}\\ e^{-i\theta}&-g\end{pmatrix},\hat{L}_{\theta}=|g|\begin{pmatrix}0&ie^{i\theta}\\ -ie^{-i\theta}&0\end{pmatrix}$ (4)
@@
$\hat{H}_{|g|}[\rho_{s}]=\textrm{Tr}[\rho_{s}\hat{L}^{2}_{|g|}]=\frac{\epsilon}{1-|g|^{2}}+O(\epsilon^{2})$ (5a)
$\hat{H}_{\theta}[\rho_{s}]=\textrm{Tr}[\rho_{s}\hat{L}^{2}_{\theta}]=|g|^{2}\epsilon+O(\epsilon^{2})$ (5b)
@@ Appendix B Superselection rule, reference frame, and quantum estimation under SSR
$|\psi\rangle=\sum_{n=0}c_{n}|n\rangle\Longrightarrow U(\psi)|\phi\rangle=\sum_{n=0}e^{in\phi}c_{n}|n\rangle$ (6)
@@
$\mathcal{U}(\rho)$ $:=\int_{0}^{2\pi}U(\psi)|\phi\rangle\langle\phi|U(\phi)^{\dagger}$ (7)
$=\sum_{n,n^{\prime}=0}\int_{0}^{2\pi}\frac{d\phi}{2\pi}e^{i(n-n^{\prime})\phi}c_{n}c^{*}_{n}|n^{\prime}\rangle\langle n^{\prime}|$
$=\sum_{n=0}|c_{n}|^{2}|n\rangle\langle n|$
@@
$\mathcal{E}(\rho_{s})$ $=(1-\epsilon)\mathcal{E}(\rho_{s}^{(0)})+\epsilon\mathcal{E}(\rho_{s}^{(1)})+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)})$
$=(1-\epsilon)\mathbb{I}^{(0)}+\epsilon\frac{\mathbb{I}^{(1)}}{2}+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)})$ (12)
@@
$\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{|g|}[\rho_{s}^{AB}],~{}~{}\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{\theta}[\rho_{s}^{AB}]$ (14)
@@
From the additivity and Monotonicity of quantum Fisher information, we have for $\mu\in\{|g|,\theta\}$
$\mathbb{H}_{\mu}[\rho_{s}^{AB}]+\mathbb{H}_{\mu}[\rho_{a}^{AB}]\geq\mathbb{H}_{\mu}[\rho_{s}^{AB}\otimes\rho_{a}^{AB}]\geq\mathbb{H}_{\mu}[\mathcal{E}(\rho^{AB})]$ (15)
@@
$\mathcal{E}(\rho^{AB})$ $=\sum_{n,m}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})$
$=\bigoplus_{n,m}\sigma^{AB}_{n,m}$ (19)
@@
$\sigma^{AB}_{n,m}$ $=\sigma^{AB0}_{n,m}+\sigma^{AB\epsilon}_{n,m}+O(\epsilon^{2})$ (20a)
$\sigma^{AB0}_{n,m}$ $=(1-\epsilon)|f_{n,m}|^{2}|0,n_{A}\rangle\langle 0,n_{A}|\otimes|0,m_{B}\rangle\langle 0,m_{B}|$ (20b)
$\sigma^{AB\epsilon}_{n,m}$ $=\frac{\epsilon}{2}\begin{pmatrix}{|f_{n,m-1}|^{2}}&{f_{n-1,m}f^{*}_{n,m-1}g}g\\ {f^{*}_{n-1,m}f_{n,m-1}g}g^{*}&{|f_{n-1,m}|^{2}},\end{pmatrix}$ (20c)
where $|0,n_{A}\rangle$ is state with 00 photon from the ancilla and $n_{A}$ photon from the source, and $\sigma^{AB\epsilon}_{n,m}$ is written in $\{|0,n_{A}\rangle|1,(m-1)_{B}\rangle,|1,(n-1)_{A}\rangle|0,m_{B}\rangle\}$ basis.
Since $\sigma_{n,m}^{A,B}$ are density operator defined on different orthogonal Hilbert spaces, it is not difficult to see that:
$\mathbb{H}_{\mu}[\mathcal{E}(\rho^{AB})]=\sum_{n,m=0}\mathbb{H}_{\mu}[\sigma_{n,m}^{AB}]$ (21)
for $\mu\in\{|g|,\theta\}$, where the QFI for the individual unnormalized state is discussed below in lemma 3 and is given as:
$\mathbb{H}_{|g|}[\sigma_{n,m}^{AB}]=\frac{\epsilon}{1-|g|^{2}}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$ (22a)
$\mathbb{H}_{\theta}[\sigma_{n,m}^{AB}]=\epsilon|g|^{2}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}]$ (22b)
for $m,n\geq 1$
@@
The quantum Fisher information contains in the unnormalized state $\sigma_{nm}^{AB}$ is contributed by the $\epsilon$ term in Eq. 20c, which is given as:
$\mathbb{H}_{|g|}[\sigma_{n,m}^{AB}]$ $=\frac{\epsilon}{1-|g|^{2}}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{H}_{\theta}[\sigma_{n,m}^{AB}]$ $=\epsilon|g|^{2}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
@@
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n,m=1}^{n+m\leq N+1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\Longrightarrow$ $\mathbb{h}[\rho^{AB}_{a}]=\sum_{k=1}^{N}\sum_{n=1}^{k}\frac{2|f_{n,k-n}|^{2}|f_{n-1,k-n+1}|^{2}}{|f_{n,k-n}|^{2}+|f_{n-1,k-n+1}|^{2}}$ (24)
@@
Now, we can give a non-tight upper bound $h_{k}$ by upper bounding $\sum_{n=1}^{k}\sqrt{p_{n}^{k}p_{n-1}^{k}}$ from the lemma below, which shows $h_{k}\leq q_{k}\cos(\frac{\pi}{k+2})\quad\quad$ and thus:
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{k=1}^{N}h_{k}\leq\sum_{k=1}^{N}q_{k}\cos(\frac{\pi}{k+2})\leq\cos(\frac{\pi}{N+2})$